# Multi-Lattice Sampling of Quantum Field Theories via Neural Operator-based Flows

### v3

@@ 3 Flows parametrised by neural operators
    The architecture should be such that it’s divergence is reasonably cheap to compute, since it will be integrated over trajectories to compute the density represented by the flow (Equation 1).
@@
The output of the model is given by convolving the input with a parametric continuous kernel $K_{\theta}:(\mathbb{R}/L\mathbb{Z})^{D}\rightarrow\mathbb{R}$. This guarantees that the first requirement is satisfied, and the second one forces $K_{\theta}$ to be spherically symmetric. Regarding the last one, we take inspiration from Chen and Duvenaud [20], and use the combination of a conditioner function $h_{i}=c(\phi_{-i})$, whose output at any coordinate is independent of the same coordinate of the input, and a transformer function $f_{i}=\tau(h_{i},\phi_{i})$ function that combines the conditioning and the input. The advantage of this architecture is that its divergence is $\sum_{i}(\partial_{2}\tau)$ and thus the capacity of $c$ can be cheaply increased.
@@
Let now $\phi\in\mathbb{R}^{N\times....\times N}$ be a discretized scalar field on a lattice. The architecture then consists of the following sequence of steps, where the subscript $\theta$ denotes trainable parameters,
@@
    Use a per-node neural network $f_{\theta}$ to embed the field values, $\phi_{emb}=f_{\theta}(\phi)\in\mathbb{R}^{c\times N\times....\times N}$.
@@
    Use a neural network to parametrize $c$-many continous spherically symmetric kernels $K_{\theta}(r)$. Let then $\tilde{K}_{\theta}$ be the evaluation of the continous kernels on the lattice.
@@
    Mask out the origin of the discrete kernel, i.e. set $\tilde{K}_{\theta}[:,\textbf{0}]=0$.
@@
    Perform a the channel-wise convolution $\phi_{emb}\star\tilde{K}_{\theta}$ and denote the result by $C\in\mathbb{R}^{c\times N\times....\times N}$. Because of the previous step, $C_{i}$ is independent of $\phi_{i}$, and we will call it the conditioner [20].
@@
    Apply a per-node neural network $\tau_{\theta}$ to the concatenation $(C,\phi_{emb})$ with output $Y=\tau_{\theta}(C,\phi_{emb})\in\mathbb{R}^{T\times N\times....\times N}$.
@@
    Contract the first dimension of $Y$ with a vector of length $T$ that only depends on time.
@@
    Finally, denoting all the above steps as $i$, we set the output of the model to be $V(\phi,t)=\tfrac{1}{2}*(i(\phi,t)-i(-\phi,t))$. This enforces the $\mathbb{Z}_{2}$ symmetry of the system.
To compute the divergence of the architecture one needs the Jacobians of the per-point operations $f_{\theta}$ and $\tau_{\theta}$, $K_{\theta}$ does not have to be differentiated through.
@@
where $p$ runs over $\left\{-\frac{\lfloor{N/2}\rfloor}{L},....,\frac{0}{L},...,\frac{\lceil{(N-1)/2}\rceil}{L}\right\}^{D}$ and the prefactor $\frac{1}{\sqrt{N^{D}}}$ makes the map $\{\phi_{x}\}\leftrightarrow\{\tilde{\phi}_{p}\}$ unitary. The covariance matrix of the free theory is diagonalized in the momentum basis $\tfrac{1}{\sqrt{N^{D}}}e^{{2\pi i\langle x,p\rangle}}$ with eigenvalues
@@
To constrain the sampling to real valued fields, we sample $p$ from the hermitian symmetric subspace of real dimension $N^{D}$ of the Fourier-space of complex dimension $N^{D}$.
@@ 4.1 Multi-lattice sampling in $D=1$ dimension
We now work in $D=1$ dimensions. Strictly speaking, a one dimensional lattice does not correspond to a quantum field theory, rather it describes the trajectory of a quantum mechanical particle in a potential. Nonetheless, it’s the simplest setup in which we can experiment and serves as a good starting point. We also fix $L=4,m^{2}=-4,g=1$ and train a single model for $5000$ steps with mesh size uniformly sampled at each training step from $N=L/a\in\{4,8,16,...,128\}$. We then evaluate performance on lattices of size $N=L/a$ up to 512 by sampling from the trained model to calculate the effective sample size (Figure 3)
@@
where $w_{i}$ is the importance weight $p(\phi_{i})/q_{\theta}(\phi_{i})$. We also estimate expectation values of the magnetization and its absolute value (Figure 3)
@@
We also compare flattened samples from the model against the one-dimensional Boltzmann density of the potential $m^{2}\phi^{2}+g\phi^{4}$ (Figure 5).
Figure 3: Experiment 4.1. $ESS,\langle M\rangle,\langle|M|\rangle$ computed from $16384$ samples at different lattice sizes. The blue crosses correspond to lattice sizes that the model was trained on, while orange dots denote lattice sizes unseen by the network during training.
Figure 4: Experiment 4.1. The two-point correlation function $G(x,y)$ computed from $16384$ samples on lattices the model was trained on (left) and on lattices the model was not trained on (right). Because of the symmetries of the task the correlation function only depends on the distance $r=|x-y|$, thus the function $G(r)$ is plotted.
Figure 5: Experiment 4.1. Flattened samples (orange histogram) of the model at different lattice sizes $N=L/a$ compared to the one-dimensional Boltzmann density $e^{-m^{2}\phi^{2}-g\phi^{4}}$ (blue curve).
@@ 4.2 Multi-lattice sampling in $D=2$ dimensions
In this experiment we work with $D=2,L=6,m^{2}=-4,g=6.975$ (the smallest system of [15]). We train for $15000$ steps with $N=L/a$ uniformly sampled from $[6,7,8,...32]$ at each training step. We evaluate the trained model on lattices up to size $64\times 64$. We report the effective sample size, as well as the expected value of the observables $M$ and $|M|$ (Table 1) and the estimated correlation function at different lattice sizes (Figure 6).
@@
Table 1: Experiment 4.2. Effective sample size and expected value of the observables $M,|M|$ computed at different lattice sizes. The four rightmost columns correspond to lattice sizes the model was not trained on.
@@
Figure 6: Experiment 4.2. The two-point correlation function $G(x,y)$ of the second experiment computed from $16384$ samples on lattices the model was trained on (left) and on lattices the model was not trained on (right). Because of the symmetries of the task the correlation function only depends on the distance $r=|x-y|$, thus the function $G(r)$ is plotted.
As in the previous experiment, the model does not extrapolate well to lattice sizes much higher than those that it was trained on. It is worth noting nonetheless that performance as a function of lattice size does not drop suddenly and it is still acceptable on slightly higher lattice size that the largest training lattice. This observation motivates the following experiment.
@@ 4.3 Faster convergence on a target lattice size by pretraining on smaller ones
In this experiment we consider the target $D=2,L=12,m^{2}=-4,g=5.276,N=64$. Instead of training directly on the $N=64$ lattice, we pretrain on a sequence of smaller lattices as they are significantly cheaper to work on. We start training on a $12\times 12$ lattice for 2000 steps, after which we train on lattices of size $16\times 16,20\times 20,24\times 24,28\times 28,32\times 32,36\times 36,40\times 40,44\times 44,48\times 48,52\times 52,56\times 56,60\times 60$ for 250 training steps each. Finally, we train on the target size $64\times 64$ for 1000 steps. As a baseline, we also train the same architecture only on the target size for the same total number of steps (6000). While the performance, as measured by the effective sample size on target lattice, is comparable after training (Table 2), the training procedure that "trained through" the smaller lattices was $\sim 2.4$-times quicker to train(Figure 7). Figure 8 shows the estimated correlation function at different lattice sizes computed from model checkpoints saved right after taking the last training step on the given lattice size.
Figure 7: Experiment 4.3. $ESS$ estimated during training on $128$ samples plotted against the number of training steps (left) and training time (right). Time required to take a single step (center). All plots contain two curves, one for the model that is trained on the sequence of increasing lattice sizes (blue) and one that is only trained on the $64\times 64$ lattice (orange). We also refer the reader to Figure 9 in the appendix that shows the $ESS$ values on all lattice sizes during training.
Figure 8: Experiment 4.3. The two-point correlation function $G(x,y)$ computed from $16384$ samples with (right) and without (left) log-scaled $y$-axis. These curves are computed from model checkpoints saved right after taking the last training step on the given lattice size. Because of the symmetries of the task the correlation function only depends on the distance $r=|x-y|$, thus the function $G(r)$ is plotted.
Table 2: Experiment 4.3. ESS values on $16384$ samples from the trained model. Since training on larger lattices degrades performance on smaller ones (Figure 9), the model is evaluated directly after the last training step has been performed on a given lattice size. The final column marked with $\flat$ denotes the baseline model.
@@ 5 Conclusion
In this work we explored the idea of using operator-based normalizing flows for sampling from the $\phi^{4}$ quantum field theory. Experiments 4.1 and 4.2 showed that models trained on a collection of lattices do not generalize zero-shot to lattice sizes much larger than those of the training set. They do generalize with a reasonable performance to lattice sizes slightly larger than the ones it has been trained on. Making use of this observation, in experiment 4.3 we show that training a model on a sequence of meshes of increasing size leads to faster training compared to training directly on the target lattice size.
@@ 6 Acknowledgement
The authors were supported by the Swiss National Science Foundation under grant number CRSII5_193716 - “Robust Deep Density Models for High-Energy Particle Physics and Solar Flare Analysis (RODEM)”. We express our appreciation of the Python ecosystem [21] and its libraries that made this work possible. In particular, the experimental part of the paper relies heavily on JAX [22], Flax [23], optax [22], numpy [24], hydra [25], matplotlib [26], jupyter [27] and weights&biases [28]. We thank Samuel Klein for discussions.

### v4

@@ 3 Flows parametrised by neural operators
    The architecture should be such that its divergence is reasonably cheap to compute, since it will be integrated over trajectories to compute the density represented by the flow (Equation 2).
@@
The output of the architecture is given by convolving the input with a parametric continuous kernel $K_{\theta}:(\mathbb{R}/L\mathbb{Z})^{D}\rightarrow\mathbb{R}$. This guarantees that the first requirement is satisfied, and the second one forces $K_{\theta}$ to be spherically symmetric. Regarding the last one, we take inspiration from Chen and Duvenaud [27], and use the combination of a conditioner function $h_{i}=c(\phi_{-i})$, whose output at any coordinate is independent of the same coordinate of the input, and a transformer function $f_{i}=\tau(h_{i},\phi_{i})$ that combines the conditioning and the input. The advantage of this architecture is that its divergence is $\sum_{i}(\partial_{2}\tau)$ and thus the capacity of $c$ can be cheaply increased.
@@
Let now $\phi\in\mathbb{R}^{N\times....\times N}$ be a discretized scalar field on a lattice. The architecture then consists of the following sequence of steps, where the subscript $\theta_{i}$ denotes trainable parameters,
@@
    Use a per-node neural network $f_{\theta_{1}}$ to embed the field values, $\phi_{emb}=f_{\theta_{1}}(\phi)\in\mathbb{R}^{c\times N\times....\times N}$, where $c$ is the number of channels.
@@
    Use a neural network to parametrize $c$-many continuous spherically symmetric kernels $K_{\theta_{2}}(r)$. Let then $\tilde{K}_{\theta_{2}}$ be the evaluation of the continuous kernels on the lattice.
@@
    Mask out the origin of the discrete kernel, i.e. set $\tilde{K}_{\theta_{2}}[:,\textbf{0}]=0$. This is done to avoid the dependence of every coordinate of the output on the same coordinate of the input and thus provide an efficient computation of the divergence [27].
@@
    Perform a the channel-wise convolution $\phi_{emb}\star\tilde{K}_{\theta_{2}}$ and denote the result by $C\in\mathbb{R}^{c\times N\times....\times N}$. Because of the previous step, $C_{i}$ is independent of $\phi_{i}$, and we will call it the conditioner [27].
@@
    Apply a per-node neural network $\tau_{\theta_{3}}$ to the concatenation $(C,\phi_{emb})$ with output $Y=\tau_{\theta_{3}}(C,\phi_{emb})\in\mathbb{R}^{T\times N\times....\times N}$.
@@
    Contract the first dimension of $Y$ with a vector of length $T$ that is the output of a neural network $\kappa_{\theta_{4}}$, taking time $t$ as the only input.
@@
    Finally, denoting all the above steps as $i$, we set the output of the architecture to be $V(\phi,t)=\tfrac{1}{2}*(i(\phi,t)-i(-\phi,t))$. This enforces the $\mathbb{Z}_{2}$ symmetry of the system.
To compute the divergence of the architecture one needs the Jacobians of the per-point operations $f_{{\theta_{1}}}$ and $K_{\theta_{2}}$ does not have to be differentiated through. Figure 3 shows a sketch of the architecture.
Figure 3: Sketch of the architecture. (a) Pointwise field embedding via $f_{\theta_{1}}$, continuous convolution with kernel $\tilde{K}_{\theta_{2}}$ to aggregate local information, and combination of the pointwise field values with neighborhood information via $\tau_{\theta_{3}}$. (b) Contracting the channels with learnable time-dependent weights given by $\kappa_{\theta_{4}}(t)$. (c) Averaging of the preceding steps over the sign of $\phi$ to enforce the $\mathbb{Z}_{2}$-symmetry of the theory.
@@
where $p$ runs over $\left\{-\frac{\lfloor{N/2}\rfloor}{L},....,\frac{0}{L},...,\frac{\lceil{(N-1)/2}\rceil}{L}\right\}^{D}$ and the prefactor $\frac{1}{\sqrt{N^{D}}}$ makes the map $\{\phi_{x}\}\leftrightarrow\{\tilde{\phi}_{p}\}$ unitary. Since the Boltzmann distribution of the free theory is a gaussian, we need to understand the structure of its covariance matrix to generate samples from it. This covariance matrix is diagonalized in the momentum basis $\tfrac{1}{\sqrt{N^{D}}}e^{{2\pi i\langle x,p\rangle}}$ with eigenvalues
@@
We can thus sample configurations by sampling the components independently in momentum space, and transform the samples back to position space with an inverse Fourier-transform. Moreover, we need to constrain the sampling to real valued fields in position space, therefore the sampling in momentum space must be constrained to the hermitian symmetric subspace of real dimension $N^{D}$ of the full momentum-space of complex dimension $N^{D}$.
@@ 4.1 Multi-lattice sampling in $D=1$ dimension
We now work in $D=1$ dimensions. Strictly speaking, a one dimensional lattice does not correspond to a quantum field theory, rather it describes the trajectory of a quantum mechanical particle in a potential [20]. Nonetheless, it’s the simplest setup in which we can experiment and serves as a good starting point. We also fix $L=4,m^{2}=-4,g=1$ and train a single model for $5000$ steps with mesh size uniformly sampled at each training step from $N=L/a\in\{4,8,16,...,128\}$. We then evaluate performance on lattices of size $N=L/a$ up to 512 by sampling from the trained model to calculate the effective sample size (Figure 4)
@@
where $w_{i}$ is the importance weight $p(\phi_{i})/q_{\theta}(\phi_{i})$. Moreover, we estimate the expectation values of the magnetization and its absolute value (Figure 4)
@@
We also compare flattened samples from the model against the one-dimensional Boltzmann density of the potential $m^{2}\phi^{2}+g\phi^{4}$ (Figure 6). All reported observables have been reweighted from the model to the target density. The observables indicate that model generalizes well to lattice sizes $4\leq N\leq 128$ (interpolation), but performance drops when $N>128$ (extrapolation). Note that before for lattice sizes in the interpolation regime ($N\leq 128$), the expected values of the observables in Figures 4 and 5 display a convergence to the underlying continuum limit.
Figure 4: Particle in double well potential (§4.1). $ESS,\langle M\rangle,\langle|M|\rangle$ computed from $16384$ samples at different lattice sizes. The blue crosses correspond to lattice sizes that the model was trained on, while orange dots denote lattice sizes unseen by the network during training. Note that the absolute magnetization converges to a value of $1.30$ as the lattice size is increased to $16$ and stabilizes at this value at larger lattice sizes.
Figure 5: Particle in double well potential (§4.1). The two-point correlation function $G(x,y)$ computed from $16384$ samples on lattices the model was trained on (left) and on lattices the model was not trained on (right). Because of the symmetries of the task the correlation function only depends on the distance $r=|x-y|$, thus the function $G(r)$ is plotted. Similarly to the absolute value of the magnetization, $G(r)$ approaches the continuum limit as the model is evaluated at increasing lattice sizes.
Figure 6: Particle in double well potential (§4.1). Flattened samples (orange histogram) of the model at different lattice sizes $N=L/a$ compared to the one-dimensional Boltzmann density $e^{-m^{2}\phi^{2}-g\phi^{4}}$ (blue curve).
@@ 4.2 Multi-lattice sampling in $D=2$ dimensions
In this experiment we work with $D=2,L=6,m^{2}=-4,g=6.975$ (the smallest system of [7]). We train a model for $15000$ steps with $N=L/a$ uniformly sampled from $[6,7,8,...32]$ at each training step. We evaluate the trained model on lattices up to size $64\times 64$. We report the effective sample size, as well as the expected value of the observables $M$ and $|M|$ (Table 1) and the estimated correlation function at different lattice sizes (Figure 7).
@@
Table 1: 2-dimensional sampling at different lattice sizes (§4.2). Effective sample size and expected value of the observables $M,|M|$ computed at different lattice sizes. The four rightmost columns correspond to lattice sizes the model was not trained on.
@@
Figure 7: 2-dimensional sampling at different lattice sizes (§4.2). The two-point correlation function $G(x,y)$ of the second experiment computed from $16384$ samples on lattices the model was trained on (left) and on lattices the model was not trained on (right). Because of the symmetries of the task the correlation function only depends on the distance $r=|x-y|$, thus the function $G(r)$ is plotted. Like in the previous experiment, $G(r)$ approaches the continuum limit as the model is evaluated at increasing lattice sizes.
As in the previous experiment, the model does not extrapolate well to lattices much larger than those that it was trained on. It is worth noting that performance as a function of lattice size does not drop suddenly and it is still acceptable on lattices slightly larger than the largest training lattice. This observation motivates the following experiment.
@@ 4.3 Faster convergence on a target lattice size by pretraining on smaller ones
In this experiment we consider a target with the following parameters: $D=2,L=12,m^{2}=-4,g=5.276,N=64$. Instead of training directly on the $N=64$ lattice, we pretrain on a sequence of smaller lattices as they are significantly cheaper to work on. We start training on a $12\times 12$ lattice for 2000 steps, after which we train on lattices of size $16\times 16,20\times 20,24\times 24,28\times 28,32\times 32,36\times 36,40\times 40,44\times 44,48\times 48,52\times 52,56\times 56,60\times 60$ for 250 training steps each. Finally, we train on the target size $64\times 64$ for 1000 steps. As a baseline, we also train the same architecture only on the target size for the same total number of steps (6000). While the performance, as measured by the effective sample size on target lattice, is comparable after training (Table 2), the training procedure that "trained through" the smaller lattices was $\sim 2.4$-times quicker to train(Figure 8). Figure 9 shows the estimated correlation function at different lattice sizes computed from model checkpoints saved right after taking the last training step on the given lattice size.
Figure 8: Pretraining on smaller lattices (§4.3). $ESS$ estimated during training on $128$ samples plotted against the number of training steps (left) and training time (right). Time required to take a single step (center). All plots contain two curves, one for the model that is trained on the sequence of increasing lattice sizes (blue) and one that is only trained on the $64\times 64$ lattice (orange). We also refer the reader to Figure 10 in the appendix that shows the $ESS$ values on all lattice sizes during training.
Figure 9: Pretraining on smaller lattices (§4.3). The two-point correlation function $G(x,y)$ computed from $16384$ samples. These curves are computed from model checkpoints saved right after taking the last training step on the given lattice size. Because of the symmetries of the task the correlation function only depends on the distance $r=|x-y|$, thus the function $G(r)$ is plotted.
Table 2: Pretraining on smaller lattices (§4.3). ESS values on $16384$ samples from the trained model. Since training on larger lattices degrades performance on smaller ones (Figure 10), the model is evaluated directly after the last training step has been performed on a given lattice size. The final column marked with $\flat$ denotes the baseline model.
@@ 5 Conclusion
In this work we explored the idea of using a neural operator-based normalizing flows for sampling from the $\phi^{4}$ quantum field theory. The main advantage of this approach is that the operator-based parametrization captures a continuous representation of the task, just like a continuous field theory underlies the lattice field theory. This in particular means that a single model can be evaluated at different lattice resolutions, and can be used to approximate the continuum limit. Experiments 4.1 and 4.2 showed that models trained on a collection of lattices do not generalize zero-shot to lattice sizes much larger than those of the training set. They do generalize with a reasonable performance to lattice sizes slightly larger than the ones it has been trained on. Making use of this observation, in experiment 4.3 we show that training a model on a sequence of meshes of increasing size leads to faster training compared to training directly on the target lattice size.
@@ 6 Acknowledgement
The authors were supported by the Swiss National Science Foundation under grant number CRSII5_193716 - “Robust Deep Density Models for High-Energy Particle Physics and Solar Flare Analysis (RODEM)”. We thank Samuel Klein for discussions.